# QPanda3: A High-Performance Software-Hardware Collaborative Framework for Large-Scale Quantum-Classical Computing Integration

### v1

@@ Abstract
QPanda3 is a high-performance quantum programming framework that enhances quantum computing efficiency through optimized circuit compilation, an advanced instruction stream format (OriginBIS), and hardware-aware execution strategies. These engineering optimizations significantly improve both processing speed and system performance, addressing key challenges in the NISQ era. A core innovation, OriginBIS, accelerates encoding speeds by up to 86.9× compared to OpenQASM 2.0, while decoding is 35.6× faster, leading to more efficient data handling, reduced memory overhead, and improved communication efficiency. This directly enhances the execution of quantum circuits, making large-scale quantum simulations more feasible. Comprehensive benchmarking demonstrates QPanda3’s superior performance: quantum circuit construction is 20.7× faster, execution speeds improve by 3.4×, and transpilation efficiency increases by 14.97× over Qiskit. Notably, in compiling a 118-qubit W-state circuit on a 2D-grid topology, QPanda3 achieves an unprecedented 869.9× speedup, underscoring its ability to handle complex quantum workloads at scale. By combining high-speed quantum processing with a modular and extensible software architecture, QPanda3 provides a practical bridge between today’s NISQ devices and future fault-tolerant quantum computing. It facilitates real-world applications in financial modeling, materials science, and combinatorial optimization, while its robust and scalable design supports industrial adoption and cloud-based deployment.
@@
Quantum Computing, Software-Hardware Collaborative, High-Performance, Quantum Circuit Compilation, Intermediate Representation
@@ I Introduction
Quantum computing, as a revolutionary computational paradigm, has garnered widespread attention from both academia and industry due to its exceptional parallel computing capabilities and potential to solve problems intractable for classical computers. In recent years, significant advancements in quantum hardware have propelled quantum computing from theoretical exploration to practical applications. However, to fully harness its potential, robust software frameworks are essential to bridge the gap between users and quantum hardware. The effectiveness of quantum computing software frameworks depends on several key factors, including user accessibility, computational efficiency, and seamless hardware integration. Against this backdrop, QPanda3 [1], a modern quantum programming framework, has undergone extensive optimization to enhance usability, quantum circuit compilation efficiency, and program transmission performance, thereby driving quantum computing toward broader applications.
@@
Currently, quantum computing is undergoing a transition from theoretical research to engineering practice. Despite the challenges faced by existing quantum hardware, such as noise, decoherence, and limited qubit scalability, engineering optimizations are steadily improving its usability, enabling Noisy Intermediate-Scale Quantum (NISQ) devices to demonstrate computational advantages over classical computers for specific problems. This engineering evolution not only determines the practical applications of NISQ devices but also influences the feasibility of large-scale fault-tolerant quantum computing in the future. Thus, enhancing the usability, stability, and execution efficiency of quantum computing has become a central challenge in its research and application. Key approaches to addressing these challenges include compilation optimization, instruction stream design, and software-hardware co-optimization.
@@
As a high-performance quantum programming framework, QPanda3 adopts an engineering-driven approach to enhance the practical usability of quantum computing. Its core objectives include:
  • Enhancing the computational capabilities of existing quantum devices—Through efficient compilation techniques, optimized qubit mapping strategies, and quantum gate compression algorithms, QPanda3 enables NISQ devices to execute target algorithms more efficiently, achieving superior computational performance for specific tasks.
  • Optimizing software-hardware co-design in quantum computing—Unlike classical computing, quantum computing requires software to have deep awareness of hardware characteristics. QPanda3 employs adaptive qubit mapping and hardware-aware compilation strategies to ensure optimal execution across different quantum processors, thereby improving overall computational fidelity.
  • Advancing quantum computational advantages—Although some NISQ devices have demonstrated computational superiority for certain problems, achieving large-scale quantum computing necessitates further engineering breakthroughs. QPanda3 adopts a modular architecture, allowing seamless integration with future fault-tolerant quantum computing systems, thereby reducing the transition cost from NISQ to next-generation quantum technologies.
The intermediate representation (IR) of quantum programs serves as a crucial bridge between high-level quantum programming languages and low-level hardware instructions, influencing program readability, portability, and execution efficiency. However, traditional quantum IR solutions often suffer from poor readability and low transmission efficiency, making it difficult for researchers to understand the mapping between high-level algorithms and low-level hardware execution. Furthermore, the lack of a unified IR standard among mainstream quantum programming languages hinders cross-platform algorithm benchmarking and experimental comparisons.
To address these issues, QPanda3 introduces a novel IR framework featuring OriginIR and an efficient binary instruction set OriginBIS to enhance program readability, transmission efficiency, and cross-platform compatibility. In classical-quantum hybrid computing and multi-device collaborative computing scenarios, existing IR solutions are typically optimized for specific hardware architectures, limiting their portability across different quantum platforms. By optimizing its IR design, QPanda3 enables efficient migration across diverse quantum platforms, thereby improving task allocation flexibility and execution efficiency.
Efficient quantum circuit compilation is crucial for transforming abstract quantum algorithms into hardware-executable instructions, encompassing quantum gate decomposition, qubit mapping, and circuit optimization. Building upon its predecessor, QPanda2, QPanda3 introduces multiple optimizations that significantly enhance compilation efficiency. Benchmarking results indicate that QPanda3 surpasses QPanda2 and other mainstream quantum computing frameworks in circuit compilation speed, dramatically improving quantum algorithm execution efficiency. In the NISQ era, where hardware constraints are prevalent, optimizing compilation efficiency is critical for maximizing computational utility.
As a quantum programming framework focused on software-hardware co-optimization, QPanda3 excels in cloud-based quantum computing environments. Given the scarcity and high cost of quantum computing resources, centralized cloud deployment is often adopted to enhance accessibility and resource utilization. Additionally, cloud-based quantum computing facilitates the establishment of standardized software-hardware environments, allowing users to conduct experiments and development without direct access to quantum hardware. Moreover, the cloud computing model fosters ecosystem collaboration in quantum computing, simplifying resource sharing and algorithm optimization.
To achieve optimal quantum execution, a comprehensive understanding of quantum hardware characteristics—such as qubit connectivity, error rates, and coherence times—is essential. QPanda3 integrates device-aware compilation techniques to minimize circuit depth and gate operations, thereby improving computational efficiency. Furthermore, QPanda3 provides device-specific quantum program analysis tools to help users assess execution performance, identify potential bottlenecks, and further optimize quantum programs. Through software-hardware co-optimization, QPanda3 enables quantum programs to adapt to different hardware architectures, achieving more efficient computation.
Experimental results demonstrate that QPanda3 significantly outperforms existing quantum programming frameworks across multiple performance metrics. For example:
  • Compilation efficiency: Experiments on Benchpress[2], in a fully connected topology, QPanda3’s compilation speed is, on average, 9.46× faster than Qiskit[3], with peak acceleration reaching 77.76×. In a square topology, its average compilation speed surpasses Qiskit[3] by 24.6×, with a peak of 869.9×. For heavy-hexagon topologies, QPanda3 achieves an average 15.1× speedup, with a peak of 332.0×. In linear topology, the average speedup is 10.71×, with peak acceleration reaching 868.7×.
  • IR transmission efficiency: The encoding speed of OriginBIS is 18.7× faster than OriginIR and 86.9× faster than OpenQASM 2.0 [4]. The decoding speed is 18.9× faster than OriginIR and 35.6× faster than OpenQASM 2.0.
The content of this paper is organized as follows: Firstly, we review the related work on quantum computing compilation frameworks. Secondly, we provide an overview of the overall architecture and core features of QPanda3. Then, we successively introduce the representation method and transmission mechanism of quantum programs, device-based quantum circuit optimization and compilation strategies, and device-based quantum program analysis, delving into the three key components of QPanda3. Finally, we validate the performance advantages of QPanda3 through experiments.
@@ II-A Intermediate Representation of Quantum Programs
Intermediate representations of quantum programs that resemble high-level programming languages offer good user readability. However, such representations tend to be relatively complex. For example, with the increasing complexity of quantum computing systems and application demands, IBM has continuously extended QASM to develop OpenQASM 2.0 and OpenQASM 3.0. As widely adopted versions, OpenQASM 2.0[4] and OpenQASM 3.0[6] not only extend quantum gate operations and qubit management but also introduce more complex syntactic structures and functional features, such as classical control flow and modular programming support. They are more akin to the style of the classical programming language C compared to the original QASM. F-QASM[7] extends QASM with feedback instructions, enhancing the efficiency of implementing measurement-based branch and loop statements. Similar intermediate representations include Scaffold[8], QCL[9], Quipper[10], ProjectQ[11], and QIR[12]. These types of intermediate representations for quantum programs sacrifice portability due to their complexity, hindering comparative experimental research by researchers.
@@ II-B Compilation, Optimization, Qubit Mapping, and Routing of Quantum Circuits
Quantum computing software such as Qiskit and QASMTrans[20] provides modern support for quantum circuit compilation. Paulihedral[21] is a compiler designed specifically for the VQE algorithm. Similarly, application-oriented compilers also include Twoqan[22], which is dedicated to QAOA circuits. CaQR focuses on the generation of dynamic circuits[23], is a compiler specifically for pulse-based devices. AutoComm[24] and QuComm[25], on the other hand, provide support for distributed devices.
@@ III Overview of QPanda3
QPanda3 is an advanced quantum programming framework that provides comprehensive support for users in quantum computing across both software and hardware aspects. As illustrated in Figure 2, QPanda3 not only offers abstractions for quantum programs and computing devices but also provides multiple related components and tools.
III-A Oriented Towards Quantum Computing
The core of quantum computing software lies in quantum programs, for which QPanda3 provides multi-level abstractions. Specifically, QPanda3 describes quantum programs using a quantum gate-based circuit model, while distinguishing between circuits without measurements and those with measurements to emphasize the difference between measurement operations and quantum gate operations. This distinction is directly reflected in QProg and QCircuit, which are part of the high-level programming language mechanism provided by QPanda3. The high-level programming language mechanism of QPanda3 uses instruction sequences within Python scripts to represent quantum programs. Users can design such quantum programs using Python and the APIs provided by QPanda3. Although high-level programming languages like Python are more user-friendly, they also present issues such as complex mapping relationships with hardware machine instructions and low transmission efficiency. QPanda3 offers a streamlined and efficiently transmittable intermediate representation of quantum programs. Furthermore, QPanda3 can compile these programs into a machine-level instruction stream format for quantum programs.
QPanda3 abstracts quantum computing hardware to meet users’ needs for executing quantum programs and conducting hardware-software co-design. In terms of computing resources, QPanda3 abstracts real quantum processors, quantum simulators, and classical processors, thereby facilitating users to efficiently utilize these resources to complete computational tasks. Regarding the physical attributes of quantum systems, QPanda3 provides noise simulation and Hamiltonian simulation, supporting users in conducting related experimental research.
@@ III-A Oriented Towards Quantum Computing
Centered around quantum programs, QPanda3 offers a rich set of components. A significant portion of these components is designed to meet the needs of hardware-software co-design, including modules specifically for quantum circuit optimization and compilation, modules tailored for device-based quantum program analysis, and application interfaces that facilitate local-cloud collaboration. Other components provide various supports for quantum programs, primarily including a quantum program translation module to ensure program portability, a quantum program visualization module to enhance design efficiency, and a variational quantum circuit module for batch generation of circuits and post-processing.
Figure 2: Framework
QPanda3 aims to fully satisfy users’ requirements for quantum computing-related work and therefore provides numerous useful tools. These tools pertain to quantum information, operators, and testing convenience. The tools related to quantum information offer multiple representations of quantum states and quantum channels, along with several information analysis utilities. The tools concerning operators provide corresponding transformation and simulation tools. The testing tools include interfaces for randomly generating quantum circuits, which can produce a large number of random circuits based on different configurations, thereby meeting users’ diverse testing needs.
III-B From QPanda2
QPanda3 fully leverages the successful experience of QPanda2 in terms of design and functionality. It continues the support for the Python language, enabling researchers and developers to quickly implement quantum algorithms using Python’s concise syntax and rich ecosystem. This design not only lowers the threshold for quantum programming but also enhances development efficiency, making it particularly suitable for beginners in quantum computing and cross-disciplinary researchers. QPanda3 introduces OriginIR as an efficient intermediate representation for quantum programs and provides conversion tools between OriginIR and OpenQASM. This design facilitates seamless migration of quantum programs across different frameworks and hardware platforms, while also providing convenience for the optimization and analysis of quantum circuits. QPanda3 supports multiple types of quantum simulators (such as full amplitude simulators and partial amplitude simulators), quantum noise simulation, Hamiltonian simulation, and management of classical registers. This comprehensive hardware abstraction capability allows researchers to verify and optimize quantum algorithms in various scenarios. To reduce the learning curve for users, QPanda3 retains the quantum circuit construction methods consistent with QPanda2. Users can quickly get started with familiar interfaces and syntax, while enjoying the performance improvements and functional extensions offered by the new framework.
III-C Enhanced Performance

### v2

@@ Abstract
In emerging quantum–classical integration applications, the classical time cost—especially from compilation and protocol-level communication—often exceeds the execution time of quantum circuits themselves, posing a severe bottleneck to practical deployment. To overcome these limitations, QPanda3 has been extensively optimized as a high-performance quantum programming framework tailored for the demands of the NISQ era and quantum–classical hybrid workflows. It features optimized circuit compilation, a custom binary instruction stream (OriginBIS), and hardware-aware execution strategies to significantly reduce latency and communication overhead. OriginBIS achieves up to 86.9$\times$ faster encoding and 35.6$\times$ faster decoding than OpenQASM 2.0, addressing critical bottlenecks in hybrid quantum systems. Benchmarks show 10.7$\times$ compilation speedup and up to 597$\times$ acceleration in compiling large-scale circuits (e.g., a 118-qubit W-state) compared to Qiskit. n high-performance simulation, QPanda3 excels in variational quantum algorithms, achieving up to 26$\times$ faster gradient computation than Qiskit, with minimal time-complexity growth across circuit depths. These capabilities make QPanda3 well-suited for scalable quantum algorithm development in finance, materials science, and combinatorial optimization, while supporting industrial deployment and cloud-based execution in quantum–classical hybrid computing scenarios.
@@
Quantum Computing, Software-Hardware Collaborative, High-Performance, Quantum-classical Integration, Quantum Circuit Compilation, Intermediate Representation
@@ I Introduction
Quantum computing, as a transformative computational paradigm, has attracted significant attention from both academia and industry due to its exceptional parallelism and its potential to tackle problems intractable for classical computers. In recent years, substantial advances in quantum hardware have propelled the field from theoretical exploration into practical application. However, fully realizing the potential of quantum computing requires robust software frameworks that bridge the gap between end-users and quantum hardware. The effectiveness of such frameworks hinges on several critical factors, including user accessibility, computational efficiency, and seamless hardware integration. In this context, QPanda3 [1], a modern quantum programming framework, has undergone extensive engineering optimizations to improve usability, circuit compilation efficiency, and program transmission performance—accelerating the practical deployment of quantum computing across diverse application domains.
@@
Quantum computing is currently undergoing a transition from theoretical exploration to engineering realization. Despite persistent hardware challenges—such as noise, decoherence, and limited qubit scalability—ongoing engineering advancements are steadily improving the usability of Noisy Intermediate-Scale Quantum (NISQ) devices. These improvements have enabled NISQ systems to demonstrate computational advantages over classical systems in specific problem domains. This evolution not only broadens the scope of near-term quantum applications but also lays the groundwork for scalable and fault-tolerant quantum computing in the future.
However, in contrast to the steady progress in hardware, most existing quantum programming frameworks exhibit a lack of engineering maturity, which poses significant barriers to the practical deployment of quantum–classical hybrid applications. In particular, the computational and communication overhead on the classical side often introduces substantial runtime latency, undermining the scheduling efficiency of quantum programs—especially in scenarios involving frequent quantum–classical interactions. These challenges are especially pronounced during quantum program compilation and in the implementation of quantum–classical communication protocols.
Many mainstream frameworks remain at a research-oriented stage, with limited attention paid to compilation efficiency, scalability for large-scale circuits, and deep integration with hardware architectures. For example, widely used plaintext-style quantum intermediate representations (QIRs), such as OpenQASM [2] and OriginIR [3], while offering human readability, suffer from overly redundant instruction formats and lack structured expressiveness. These limitations complicate backend optimization and hinder efficient quantum–classical co-execution, exacerbating communication latency and creating scheduling bottlenecks. Moreover, plaintext QIRs introduce additional obstacles in real-world deployments, including increased parsing overhead, inefficient transmission, and security vulnerabilities. These issues collectively restrict the development of efficient, reliable, and scalable quantum compilation and execution infrastructures.
Addressing these limitations requires a fundamental rethinking of quantum programming frameworks from an engineering perspective. This includes optimizing the compilation pipeline, designing efficient intermediate representations, and enabling high-performance execution tailored to the demands of practical quantum–classical systems.
@@
QPanda3, as a high-performance quantum programming framework, adopts an engineering-centric design philosophy aimed at enhancing the practical usability of quantum computing systems. Its core objectives include:
  • Superior Compilation Performance:
    QPanda3 employs a customized data structure tailored to the quantum compilation workflow, effectively reducing intermediate-stage information redundancy in instruction reordering, topology-constrained mapping, and gate fusion optimization. This design enables efficient and scalable compilation for quantum circuits at the scale of thousands of qubits. Benchmark results from Benchpress [4] demonstrate that, in fully connected topologies, QPanda3 achieves an average compilation speed 7.03$\times$ faster than Qiskit [5], with peak speedups up to 123.33$\times$. For square, heavy-hexagon, and linear topologies, average and peak speedups reach 15.87$\times$/597.41$\times$, 12.45$\times$/313.86$\times$, and 7.43$\times$/374.48$\times$, respectively.
  • Quantum-Classical Protocol Optimization:
    By introducing a concise and efficient binary stream format to represent quantum programs, QPanda3 significantly improves encoding and decoding speeds as well as compression ratios compared to traditional plaintext intermediate representations. Specifically, the encoding speed of OriginBIS surpasses that of OriginIR and OpenQASM 2.0 by factors of 18.7$\times$ and 86.9$\times$, respectively; decoding speed is enhanced by 18.9$\times$ and 35.6$\times$. This advancement effectively reduces communication latency in quantum hyper-converged applications.
  • Fast Simulation of Hybrid Quantum Programs:
    QPanda3 provides modular support for efficient simulation of variational quantum circuits and quantum-classical hybrid algorithms, addressing the rapid iteration and feedback-driven development characteristic of the NISQ era. Comparative experiments with Qiskit, PennyLane [6], and DeepQuantum [7] confirm that QPanda3 achieves the shortest gradient computation times and overall superior performance.
  • Quantum Software and Hardware Profiling:
    The framework incorporates a comprehensive performance analysis toolchain supporting benchmarking, calibration, and cross-platform performance comparison of both quantum software and underlying hardware, facilitating the validation and advancement of emerging quantum technologies.
The remainder of this paper is organized as follows: Section 2 reviews related work on quantum compilation frameworks. Section 3 outlines the architecture and key features of QPanda3. Section 4 presents experimental validation of QPanda3’s performance advantages. In addition, the Appendix provides a modular breakdown of QPanda3’s internal components along with illustrative code examples, enabling readers to gain a deeper understanding of its usage and extensibility in practical quantum programming scenarios.
@@ II-A Intermediate Representation of Quantum Programs
Intermediate representations of quantum programs that resemble high-level programming languages offer good user readability. However, such representations tend to be relatively complex. For example, with the increasing complexity of quantum computing systems and application demands, IBM has continuously extended QASM to develop OpenQASM 2.0[2] and OpenQASM 3.0 [9]. As widely adopted versions, OpenQASM 2.0 and OpenQASM 3.0 not only extend quantum gate operations and qubit management but also introduce more complex syntactic structures and functional features, such as classical control flow and modular programming support. They are more akin to the style of the classical programming language C compared to the original QASM. F-QASM[10] extends QASM with feedback instructions , enhancing the efficiency of implementing measurement-based branch and loop statements. Similar intermediate representations include Scaffold [11], QCL[12], Quipper [13], ProjectQ[14], and QIR[15]. These types of intermediate representations for quantum programs sacrifice portability due to their complexity, hindering comparative experimental research by researchers.
@@ II-B Compilation, Optimization, Qubit Mapping, and Routing of Quantum Circuits
Quantum computing software such as Qiskit and QASMTrans[23] provides modern support for quantum circuit compilation. Paulihedral [24] is a compiler designed specifically for the VQE algorithm. Similarly, application-oriented compilers also include Twoqan [25], which is dedicated to QAOA circuits. CaQR[26] focuses on the generation of dynamic circuits, is a compiler specifically for pulse-based devices. AutoComm[27] and QuComm [28], on the other hand, provide support for distributed devices.
@@ II-C Quantum Program Profiling
II-D Variational Quantum Circuit
Variational quantum algorithms [41] and quantum machine learning [42], as two important research fields, represent two key directions and milestones in the practical application of quantum computing. The Variational Quantum Eigensolver (VQE) has been widely applied in quantum chemistry, condensed matter physics, and optimization problem-solving. The Quantum Approximate Optimization Algorithm (QAOA) [37] can be employed for combinatorial optimization problems such as the Max-Cut problem [43], the Traveling Salesman Problem, portfolio optimization, and scheduling. The Variational Quantum Classifier [44] is a significant application of variational quantum algorithms in the field of machine learning. Meanwhile, other quantum machine learning techniques, such as quantum neural networks, may also utilize non-variational quantum algorithms. In quantum computing, most of these critical algorithms rely on parameterized quantum circuits. Therefore, parameterized quantum circuits serve as essential components for the practical implementation of quantum computing.
Variational quantum algorithms adopt the term Ansatz [45]. In algorithms such as VQE [36], QAOA, and VQC, the term Ansatz can unambiguously refer to a parameterized quantum circuit while also implying the optimization of parameters to solve a target problem. However, Ansatz models that are not based on quantum circuits also exist. QPanda2 [46] explicitly uses variational quantum circuit to refer to parameterized quantum circuits employed in variational algorithms [47]. It is worth noting that parameterized quantum circuits designed for variational algorithms can also be repurposed as general parameterized circuits and applied in quantum machine learning algorithms that do not rely on variational methods. QPanda3 retains the term *variational quantum circuit* from QPanda2, with parameterized quantum circuits at its core. It provides functionalities such as quantum circuit generation, Hamiltonian expectation value calculation, and gradient computation, thereby supporting the use of parameterized quantum circuits in both variational quantum algorithms and quantum machine learning algorithms. This differs from frameworks like Qiskit [5], PennyLane [6], and MindQuantum [48], which do not explicitly distinguish between standard quantum circuits (without tunable parameters) and parameterized quantum circuits. In contexts related to QPanda3, where no ambiguity arises, the terms Ansatz, parameterized quantum circuit, and variational quantum circuit are used interchangeably without explicit distinction.
QPanda3 is a foundational quantum programming framework designed for cross-disciplinary research. While quantum machine learning-focused frameworks such as PennyLane, MindQuantum, and pyVQNet [49] have been optimized for various aspects of machine learning, they often struggle to support quantum computing tasks in other domains. A similar limitation exists with ChemiQ [50], which provides comprehensive support for quantum chemistry but is specialized for that field. QPanda3 maintains good compatibility with pyVQNet and ChemiQ. However, unlike these frameworks-which incorporate extensive application-specific features tailored to their respective domains-QPanda3’s variational quantum circuits do not extend many application-dependent functionalities. Instead, they focus on providing efficient implementations of essential core operations. By offering more fundamental building blocks, QPanda3 enables researchers to use a single software platform to customize algorithmic implementations across different application areas, thereby facilitating cross-disciplinary research. These essential core functionalities include: Constructing parameterized quantum circuits, Obtaining parameterized quantum states for given parameter values, Reusing variational quantum circuit structures for modular programming, Computing Hamiltonian expectation values, Calculating gradients, and Supporting expressions as parameters in quantum gates. This design philosophy allows QPanda3 to serve as a versatile and adaptable tool for researchers exploring diverse quantum computing applications.
QPanda3 is a quantum programming framework designed for high-performance computing. While parameterized quantum circuits can be implemented by combining function definitions in programming languages with the general quantum circuit construction mechanisms provided by programming frameworks, this approach imposes limitations on certain high-performance extensions. For instance, when computing Hamiltonian expectation values and gradients, more efficient implementations exist on classical computers-some of which rely on direct manipulation of quantum states stored in classical memory [51]. Currently, quantum programming SDKs such as Qiskit, PennyLane, pyQuil, MindQuantum, and QPanda all offer Python-based interfaces. However, constructing parameterized circuits using Python functions and generic circuit-building mechanisms prevents fine-grained control over memory operations, introduces redundant instructions, and ultimately restricts performance improvements in variational quantum circuit functionalities. To address this, QPanda3 provides dedicated data structures and processing logic specifically optimized for variational quantum circuits, aiming to enhance the performance of variational quantum computing tasks.
@@ III Overview of QPanda3
QPanda3 is an advanced quantum programming framework that provides comprehensive support for quantum computing across both software and hardware domains. As illustrated in Figure 2, QPanda3 offers abstractions for quantum programs and computing devices, along with a wide array of supporting components and tools.
III-A Quantum Computing Orientation
At the core of quantum software lies the quantum program, for which QPanda3 provides multi-level abstractions. Specifically, QPanda3 adopts a quantum gate-based circuit model, distinguishing between circuits with and without measurements. This distinction emphasizes the differing roles of measurement operations and quantum gate operations and is directly embodied in the QProg and QCircuit structures provided by QPanda3’s high-level programming interface.
Quantum programs are expressed as instruction sequences in Python, allowing users to design quantum circuits using QPanda3’s APIs. While Python enhances usability, it introduces complexities in hardware instruction mapping and transmission inefficiency. QPanda3 addresses these issues by introducing a streamlined intermediate representation (IR) that supports efficient encoding and transmission, and by compiling programs into hardware-level instruction formats.
For hardware abstraction, QPanda3 encompasses real quantum processors, quantum simulators, and classical computing units, enabling users to leverage diverse resources. Additionally, it supports noise modeling and Hamiltonian simulation to facilitate physical-system-level experimentation.
@@ III-A Quantum Computing Orientation
Surrounding the quantum program, QPanda3 incorporates a rich set of components designed to enable hardware–software co-design. These include modules for circuit optimization and compilation, device-specific quantum program analysis, and cloud-based execution interfaces. Other auxiliary tools include quantum program translation for portability, visualization tools for circuit design, and a variational circuit module for efficient circuit generation and post-processing.
Variational quantum circuits—key elements of quantum-classical hybrid computing—are also natively supported in QPanda3.
Figure 2: Architecture of the QPanda3 Framework
To further support quantum development, QPanda3 provides tools for quantum state and channel representation, operator simulation, and testing. These include utilities for state transformation, information-theoretic analysis, and random circuit generation, catering to diverse testing and benchmarking needs.
III-B Advancements from QPanda2
QPanda3 builds on the successful design principles of QPanda2. It continues to support Python, offering a concise syntax and rich ecosystem, thereby lowering the entry barrier and accelerating development—particularly for interdisciplinary researchers.
A key innovation in QPanda3 is the introduction of OriginIR, an efficient intermediate representation, along with conversion tools for interoperability with OpenQASM. This enables seamless migration across platforms and supports program optimization and analysis. The framework also supports multiple simulator types, noise and Hamiltonian simulations, and classical register management, thereby offering a robust hardware abstraction layer.
To ease the transition from QPanda2, QPanda3 maintains compatibility with its circuit construction paradigm, allowing users to reuse familiar interfaces while benefiting from enhanced performance and functionality.
III-C End-to-End Software-Hardware Co-Design in QPanda3
To address the various challenges encountered during the execution of quantum algorithms on real hardware, QPanda3 adopts a software-hardware co-design paradigm that spans the entire workflow including program design, transmission, compilation, analysis, and execution. Figure 3 illustrates the framework’s full-stack integration process in both local and cloud environments.
Figure 3: Workflow of QPanda3 emphasizing software-hardware co-design. The framework realizes an integrated closed-loop encompassing program construction, cloud-based transmission, device-oriented compilation, and adaptive performance analysis.
III-C1 Cloud Execution and Cross-Device Program Transmission
With the increasing adoption of the “local design, cloud execution” paradigm in quantum computing, QPanda3 offers the following capabilities to support this scenario:
  • Employs a compact intermediate representation (IR) that supports efficient, compressed, and encrypted transmission;
  • Provides APIs and CLI tools to enable seamless interoperability with cloud backend systems;
  • Supports device introspection interfaces capable of retrieving target quantum device topology, calibration data, and error rates.
These capabilities empower users to efficiently deploy quantum programs across heterogeneous platforms, thereby enabling truly cloud-native quantum computing operations.
III-C2 Hardware-Oriented Collaborative Optimization Compilation Mechanism
QPanda3 utilizes a dynamic feedback-driven compilation architecture that constructs quantitative evaluation models based on real-time hardware performance metrics (e.g., gate fidelity, decoherence times, crosstalk noise) to drive multi-level compilation optimizations. The core innovation lies in establishing a closed-loop optimization system of compilation–execution–feedback:
Hardware-Aware Real-Time Tuning Mechanism
  • Dynamically monitors chip topology, gate error rates, and other parameters at runtime to generate adaptive mapping strategies;
  • Supports dynamic instruction set redirection for heterogeneous quantum chips (e.g., superconducting, trapped ion).
Performance-Driven Layered Compilation
  • Physical Layer: Gate decomposition strategies dynamically adjusted based on chip T1/T2 parameters, optimizing parallel gate scheduling;
  • Logical Layer: Incorporates backend feedback on chip information to optimize layout and routing;
  • System Layer: Orchestrates pipeline scheduling of hybrid classical-quantum instructions to enhance quantum resource utilization.
Heterogeneous Computing Adaptation Framework
  • Constructs a unified intermediate representation layer (QIR) supporting cross-platform instruction translation;
  • Continuously updates a chip characteristics database (QCDB) including gate sets, topology, and noise models;
  • Employs adaptive backend code generators optimized respectively for control timing on superconducting and trapped ion chips.
III-C3 Quantum Program Analysis and Optimization
QPanda3 provides a comprehensive quantum program analysis toolchain supporting multi-dimensional program characteristic evaluation and optimization from the logical layer down to the physical layer. The main analysis functions include:
Program Structure Analysis
  • Quantum gate-level dependency analysis;
  • Identification of parallelism and critical paths;
  • Control flow analysis of quantum-classical hybrid programs.
Performance Evaluation
  • Supports fidelity evaluation methods such as cross-entropy benchmarking (XEB) and randomized benchmarking;
  • Complexity analysis including quantum circuit depth and gate count.
Resource Bottleneck Analysis
  • Statistics on gate-level parallelism and qubit utilization;
  • Detection of routing conflicts under qubit topology constraints;
  • Pulse-level scheduling and timing analysis.
III-D Enhanced Performance